Arrange the following gases in order of decreasing density at STP:

HCl, HBr, Ne, O2
Answer #1

Answer

HBr > HCl > O2 > Ne

Explanation

Ideal gas equation is

PV = nRT

n/V = P/RT

n/V = molarity of gas

at same temperature and pressure molarity of gases would be same

mass = mole × molar mass

as molar mass of a gas higher mass per liter would be higher , so density of the gas would be higher.